The sum of two numbers is 33. Their difference is 9. What are the two numbers?
Answer by mathslover(132)

Let x and y be the 2 numbers
then we have
x + y =33 ........(1) (Sum is 33 )
and
x - y =9 ........(2) (difference is 9 )
adding 1 and 2
2x = 42
x= 21
Substituting the value of x in 1
21 + y = 33
y = 33-21
y= 12
Numbers are 21 and 12